"Find the intercepts for the equation"
4x-y=4
x-intercept:
y-intercept:

Answer by rfer(16322)

x int
4x-0=4
x=4/4
x=1
(1,0)
------------
y int
-y=-4x+4
y=4x-4
(0,-4)